OpenStudy (anonymous):

Draw the graph of the circle (x+1)^2+(y+4)^2=4

OpenStudy (sweetburger):

do you know how to find your center points?

OpenStudy (anonymous):

the center of \[(x-h)^2+(y-k)^2=r^2\] is \((h,k)\) and the radius is \(r\)

OpenStudy (anonymous):

my center points are (-1,-4) correct? @satellite73

OpenStudy (anonymous):

@sweetburger

mathslover (mathslover):

Yeah , that's right @xForeverASaintx

OpenStudy (sweetburger):

@xForeverASaintx yes you would be correct

OpenStudy (anonymous):

ok so then how do i draw it?

OpenStudy (sweetburger):

you have your center so make a dot on the graph there now look for your radius

OpenStudy (anonymous):

my radius is 2

OpenStudy (anonymous):

yes to both

OpenStudy (anonymous):

but she wants me to draw it and i have no graph i literally have to draw it with a digital pen

OpenStudy (anonymous):

good luck put a dot at \((-1,-4)\) and draw some circle about that dot with radius 2

OpenStudy (anonymous):

ohk sounds simple enough i guess

OpenStudy (anonymous):

you can make sure it goes through some points, like \((-1.-2)\) at the top and \((-1,-6)\) at the bottom who is this "she"?

OpenStudy (anonymous):

my online teacher

OpenStudy (anonymous):

tell your on line teacher that the question The equation \(y^2+4y+4x^2+4x+21=0\) can be changed into a vertex equation by completing the square. Which conic section can be made from the graph of this equation? A. ellipse B. hyperbola C. circle D. parabola was stupid

OpenStudy (anonymous):

ok don't

OpenStudy (anonymous):

ha ha ill be sure to NOT let her know that. lol

OpenStudy (anonymous):

good idea why ruffle her feathers she probably didn't write the question anyway

OpenStudy (anonymous):

ha ha yea you're right im sure she didnt